A runner’s heart rate in beats per minute is given by the formula Heart rate = 65 + 53 / (4t + 1) where t is the number of minutes after the start of a cool-down period. Find the runner’s heart rate in 10 minutes. When is the runner’s heart rate equal to 68 beats per minute?
Answer by CharlesG2(834)

A runner’s heart rate in beats per minute is given by the formula Heart rate = 65 + 53 / (4t + 1) where t is the number of minutes after the start of a cool-down period. Find the runner’s heart rate in 10 minutes. When is the runner’s heart rate equal to 68 beats per minute?
Heart rate = 65 + 53 / (4t + 1) (plug in 10 minutes for t)
HR = 65 + 53/(4*10 + 1) (4*10=40)
HR = 65 + 53/41
HR = 66.292683
to find 68 beats/min 53/(4t+1) = 3 and solve for t
53/(4t+1) = 3
53=3*(4t+1)
53=12t+3
50=12t
50/12=25/6=4 1/6 minutes = t = time to get HR to equal 68 beats per minute